OpenStudy (anonymous):

first group has 25 students and 2 teachers and cost $97.50 second group has 32 students and 3 teachers and cost $127.00 what is the cost per student and teacher?

OpenStudy (anonymous):

25s + 2t = 97.50 32s + 3t = 127.00 solve via substituion or elimination

OpenStudy (anonymous):

these are the equations i tried and couldn't get them to come out right

OpenStudy (anonymous):

let me calculate

OpenStudy (anonymous):

i get s = 3.5 and t =5 ... what did you get?

OpenStudy (anonymous):

so means students cost 3.50 and teachers cost 5.00

OpenStudy (anonymous):

s=1.86 and knew that was wrong so didn't try for t

OpenStudy (anonymous):

let me draw the answer - mine is correct.

OpenStudy (anonymous):

25s +2t = 97.50 multiply by -3 32s +3t = 127 multiply by 2 -75s + -6t = -292.50 64s + 6t = 254.00 now combine, -11s = -38.50 s = 3.50

OpenStudy (anonymous):

25s + 2t = 97.50 25 (3.50) + 2t = 97.50 87.50 - 97.50 = -2t -10 = -2t t = 5.00

OpenStudy (anonymous):

oh you did by elimination and i was trying substitution and couldn't get it to come out right thank you

OpenStudy (anonymous):

substition would have worked. do you want me to do that way also - then you can see the step where an error was made - just let me know

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

25s + 2t = 97.50 32s + 3t = 127 ** 3t = 127 - 32s t = 127/3 - 32/3s

OpenStudy (anonymous):